OpenStudy (anonymous):

Does anyone know how to do this? For this adventure, you and world renowned Professor Sherlock McMerlock are traveling to the Lost Island of Laplaya. When the boat arrives on the island shore, you and Professor McMerlock disembark on your adventure. You trudge through the jungles and arrive at three impressively large doors. About eye level on each door is an intricately carved keyhole. Directions are scratched into the wood above each keyhole.

OpenStudy (anonymous):

To open these doors, you must match the number and type of solutions for the following two functions in standard form. f(x) = x^2 + 6x - 16 g(x) = x^2 + 6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrational solutions Two real rational solutions After matching these functions, explain to Professor McMerlock how you know these functions meet each condition. Remember, he is a professor, so use complete sentences.

OpenStudy (shadowlegendx):

SHERLOCK HOLMES ;O